(1 point) Consider the linear system a. Find the eigenvalues and eigenvectors for the coefficient matrix. b. Find the real-va Please answer all parts and box answers
Answer #1

S -3 07-7-) –12](Mati (-5) = 0 For iguales les o I + 4+1=0 = 1 = for dei ( ) = (*) = (3-1), +2V2 =0 -2 ; = 3-1 -2. V. (3-0) ļbut you (o)=900)*10) 7 -261-1 174 E 7: Come 366E-5 -G = -52 67 - 2 sint i ý = 1/-2 cost 3 sint cost I cost & Sint / 5 = (sta